Find the probability of getting at least one girl when a couple has five children. Assume that boys and girls are equally likely and that the sex of any child is independent of the others.
Answer by stanbon(48568)

Find the probability of getting at least one girl when a couple has five children. Assume that boys and girls are equally likely and that the sex of any child is independent of the others.
-----------------
The easiest way to figure this is to find P(no girls in 5 births) = (1/2)^5
= 1/32
--------------
The P(at least one girl in 5 births) = 1 - 1/32 = 31/32
================
Cheers,
Stan H.